how do you go about solving log(x-5)+log(x-2)=1?
Answer by stanbon(26278)

how do you go about solving
log(x-5)+log(x-2)=1
log[(x-5)(x-2)] = 1
(x-5)(x-2) = 10
x^2-7x+10 = 10
x^2-7x = 0
x(x-7) = 0
Positive solution:
x = 7
--------------
Cheers,
Stan H.